OpenStudy (anonymous):

Vectors/line integral help please :) Find the work done by the force field: F=(x+y)i + (x-z)j +(z-y)k in moving an object from (1,0,-1) to (0,-2,3) along a smooth path. Have tried this over and over and cant seem to get the right answer.

OpenStudy (alexwee123):

how much does the object weigh?

OpenStudy (anonymous):

Not sure, it wasnt given in the question. The context of the question is line and surface integrals where the answer is of form \[W=\int\limits_{}^{}f.dr\] where f and r are vectors.

OpenStudy (anonymous):

use the following parameters x = 1-t y = -2*t z = -1+4*t they were gotten by (1-t)*r0+t*r1 Gives you the parametrization of a line segment.

OpenStudy (anonymous):

1/2 :)

OpenStudy (binary3i):

\[dr = dxi +dyj +dzk\] \[f=(x+y)i +(x-z)j +(z-y)k\] \[f.dr=(x+y)dx +(x-z)dy +(z-y)dz\]

OpenStudy (anonymous):

using the paremeters find the derivitives of each one. put them in x' i +y' j + z' k form then plug in into your vector function like this (1-t)*i+(-2*t)*j+(-1+4*t)*k take dotproduct and get -7+15*t and intigrate that from 0 to 1 wrt

OpenStudy (anonymous):

I made a mistake plugging back in

OpenStudy (ajprincess):

@binary3i can u tell me ur final answer plzzz. I too did in this way. Wanna check the answer.

OpenStudy (anonymous):

-9+37*t is when you plug back to F

OpenStudy (anonymous):

19/2 is my new answer

OpenStudy (ajprincess):

I thnk it shud b x^2/2 and z^2/2. Isnt t?

OpenStudy (anonymous):

It looks wrong to me

OpenStudy (anonymous):

binary's

OpenStudy (anonymous):

there has to be a derivitives somewhere.

OpenStudy (binary3i):

yeah i made a mistake, their i fought to divided the X^2 and Z^2 bye 2

OpenStudy (ajprincess):

Nw what do u get @binary3i

OpenStudy (anonymous):

19/2 I am 99.99999% sure of it :)

OpenStudy (ajprincess):

Can u explain ur method a bit in details if u dnt mind @timo86m.

OpenStudy (ajprincess):

plzzz

OpenStudy (binary3i):

its 31/2

OpenStudy (anonymous):

use the following parameters x = 1-t y = -2*t z = -1+4*t they were gotten by (1-t)*r0+t*r1 r0 is intial r1 is end point Gives you the parametrization of a line segment. ---------- now do this r(t)= 1- t i + -2t j + -1+4t z r'(t)=-i-2*j+4*k F(x,y,z)= (x+y)i + (x-z)j +(z-y)k the original equation Use your parameters to plug into z y and z as in ((1-t) + ( -2*t)) i +.... and so on take dot product of F(x,y,z) dot r'(t) I got -7+15*t then just intigrat that from 0 to 1 with respect to x :)

OpenStudy (anonymous):

x y and z not z y and z :P

OpenStudy (ajprincess):

I too am getting 19/2 in my method.:) Thanxxx a lot for the explanation.

OpenStudy (anonymous):

OOH yay :D

OpenStudy (anonymous):

similar problem from book

OpenStudy (binary3i):

but then whats wrong with mine?

OpenStudy (ajprincess):

Ohh no am getting both of the answers.:(

OpenStudy (anonymous):

several problems I can tell right away because the results have to be a number you taking integrals of with 2 variables will result in a variable left over. as in take integral of x + y wrx you get y left over

OpenStudy (binary3i):

\[\int\limits\limits_{}^{}f.dr= x^2/2 + 2xy -2 yz + z^2/2=I\] now do the limits on I

OpenStudy (anonymous):

show me your method later princess :) I am going to bed. i'll look later

OpenStudy (ajprincess):

gotta go nw. will show it later.:)

OpenStudy (anonymous):

OP did you find the answer?

OpenStudy (ajprincess):

nope. I dnt knw to work this out using line integrals. cos I havn't learnt t yet. I tooin binary's method bt was unabl to get the final answer.

OpenStudy (anonymous):

It is not going to work. BInaries method is wrong formula. It doesn't even have a derivitive.

OpenStudy (binary3i):

why do you want a derivative in between

OpenStudy (ajprincess):

u can use this formula ven the question is like this Find the work done by the force F=(3x^2+6yz)i+(2y+3xz)j+(1+5xyz^2)k in moving a particle from (0 0 0) to (1 1 1) (i)along the straight line x=y=z, which joins these two points.

OpenStudy (anonymous):

Here is why

OpenStudy (anonymous):

you cant just say dx you have to say dx= x'(t) dt

OpenStudy (binary3i):

only when i want them to be the function of t

OpenStudy (anonymous):

You have to be in terms of t also as I said before Another problem is this you have int(x+y) wrx from 1-0 would evaluate to -1/2-y You can tell it is instantly wrong because it is supposse to be a number with no variables.

OpenStudy (binary3i):

you see that I up there. put substitute (0,-2,3) and subtract with I substituted with (1,0,-1). u'll get it.

OpenStudy (anonymous):

Wow, thank you everyone for the response! I have had internet problems since posting but am back online :) Thanks for the help. Timo the answer is 19/2 in the back of the book- so well done. I am still having trouble understanding how to do it as my lecturer didnt cover the material necessary for your approach but will look at the examples youve posted. Thanks again all.

OpenStudy (anonymous):

See I was right :D

OpenStudy (anonymous):

use the following parameters x = 1-t y = -2*t z = -1+4*t they were gotten by (1-t)*r0+t*r1 r0 is intial r1 is end point Gives you the parametrization of a line segment. ---------- now do this r(t)= 1- t i + -2t j + -1+4t z r'(t)=-i-2*j+4*k F(x,y,z)= (x+y)i + (x-z)j +(z-y)k the original equation Use your parameters to plug into z y and z as in ((1-t) + ( -2*t)) i +.... and so on take dot product of F(x,y,z) dot r'(t) I got -7+15*t then just intigrat that from 0 to 1 with respect to x :)

OpenStudy (anonymous):

Your parametrisation is beyond me lol. we havent learnt that- is there any other approach?

OpenStudy (anonymous):

well ajppreincess reached the right answer 2 using a method I am not sure about.

OpenStudy (anonymous):

Well we learned the parametric form of line in ch12.5 or 12.6 then we used that in ch 15.2

OpenStudy (anonymous):

So (1-t)*r0+t*r1 is the general equation of how to parametrise ? im sorry im not understanding very well haha. Could you show me some steps if u have time?

OpenStudy (anonymous):

where r0 is (x0,y0,z0) initial point and r1 is just (x1,y1,z1) end point from (1,0,-1) to (0,-2,3) to get x you have x=(1-t)*(x0)+t*(x1) ... (1-t) (1)+t*(0) and you just simplify rearrange and such x=1-t

OpenStudy (anonymous):

do same for y and z

OpenStudy (anonymous):

great! i got them now. I had just never seen that equation before! very handy. Is it only for straight lines?

OpenStudy (anonymous):

Yup only for straigh line. It is for a line segment :)

OpenStudy (anonymous):

but the question didnt really state it was a line segment- how do we know it wasnt a weird curve? I am really struggling in this topic- can you recommend a really good online resource that explains line and surface integrals really clearly? Our lecturer rushed through it and the exam is soon!

OpenStudy (anonymous):

there is something called path independance If the Force field is conservative :)

OpenStudy (anonymous):

ok- if the curl F is zero?

OpenStudy (anonymous):

there is a teste for it I think if you take the cross product of it and it =0

OpenStudy (anonymous):

like walking up stairs. if you wanna get to 5th floor you waste same amnt of work aka energy If you take ladder straight up curling stairs zig zaging stair ...

OpenStudy (anonymous):

Also why did we integrate from 0 to 1?

OpenStudy (anonymous):

that is due to the paremeter formula I gave you t goes fro 0 to 1 :)

OpenStudy (anonymous):

Always?

OpenStudy (anonymous):

try it out x = 1-t y = -2*t z = -1+4*t plug in t as 0 than 1 and you should get (1,0,-1) to (0,-2,3)

OpenStudy (anonymous):

Yes always

OpenStudy (anonymous):

brilliant! thanks

OpenStudy (anonymous):

It is just part of the formula (1-t)*r0+t*r1 where 0<t<1

OpenStudy (anonymous):

great! you've been a great help thanks. can you recommend any good websites for line and surface integrals?

OpenStudy (anonymous):

Oh sorry I don know :( but I love khanacademy.org IDK if he has them. WHat he does have tho is great :)

OpenStudy (anonymous):

Oh great thanks alot!

OpenStudy (anonymous):

Ive heard that website is good- but had forgotten!

OpenStudy (anonymous):

It is very good but I dont remember seeing no line integrals there . I think I did but i was surprised on how he usually explains it well but not that time. It was so long ago tho.

OpenStudy (anonymous):

Cool. Ill dig around for something! Thanks again. Good luck with your studies.